a mixture of co and co2 is found to have a density of 1.5 gm/lit at 30  degree celsius and 730 mm .what is the composition of the mixture.

Grade:11

4 Answers

Sakshi
8 years ago
Hi..

For a mixture of CO and CO2, d = 1.50 g/litre

P = 730 / 760 atm, T = 303K

PV =w / m RT; PV w / Vm RT

730 / 760 = 1.5 / m × 0.0821 × 303 =∴ 38.85

i.e. molecular weight of mixture of CO and CO2= 38.85

Let % of mole of CO be a in mixture then

Average molecular weight = a × 28 + (100 – a) 44 / 100

38.85 = 28a + 4400 – 44a / 100

a = 32.19

Mole % of CO = 32.19

Mole % of CO2= 67.81
Regards
rutwik jadav
4 years ago
you not proprly explain it ,now let’s see my explanation that you falling into the movie genius ...P = 730  / 760  atm, T =  303KPV = w / m RT; PV w  / Vm RT730 / 760 = 1.5 / m × 0.0821 × 303 =    ∴ 38.85Average molecular weight =  a × 28 + (100 – a)  44 / 10038.85 = 28a +  4400 – 44a /  100a =  32.19Mole % of CO = 32.19Mole % of CO2= 67.81                thank you..................
Avni Sharma
4 years ago
hi friends... are you all studying for the IIT exam also? I am new here. I read your question just yesterday. I think that this question was also asked in the exam last year. I remember solving this in previous year jee question paper 2019. yEs the two answers are right friends. If you are gave the exam this year, how was it? Did you check your answers? I did that after I read that NTA realeased official answer key. You should also read take a look at this. 

blog.mypat.in/engineering-entrance-exam/jee-main-2020-paper-1-analysis-9th-january-2020
ankit singh
3 years ago
730 / 760 equals 1.5/ m which when multiplied by 0.0821 × 303 results in 38.35 which is molecular weight of the mixture of CO2 and CO. is 67.32% in the mixture. as follows.